OpenStudy (anonymous):

Which has the same value as (4 + 3i)²?

OpenStudy (luigi0210):

distribute

Parth (parthkohli):

\[(4 + 3i)^2 = 4^2 + 2(4)(3i) + (3i)^2\]

OpenStudy (anonymous):

does that mean i treat it as a variable??

Parth (parthkohli):

Yes, you can treat it as a variable. Just keep in mind that \(i^2=-1\) and that's it.

OpenStudy (anonymous):

im still confused

Parth (parthkohli):

Do you know that \((x + y)^2 = x^2 + 2xy + y^2\)?

OpenStudy (anonymous):

Yea

OpenStudy (luigi0210):

Where ever there is an i^2 sub it with a -1

Parth (parthkohli):

Using that property would suffice.\[(4 + 3i)^2 = 4^2 + 2(4)(3i) + (3i)^2\]

Parth (parthkohli):

@Luigi0210 is correct; though you want to simplify the expression I posted above.

OpenStudy (dan815):